# GiGL: Large-Scale Graph Neural Networks at Snapchat

### v1

@@ Acknowledgements.
- Zhu et al. (2021b) Zhaocheng Zhu, Zuobai Zhang, Louis-Pascal Xhonneux, and Jian Tang. 2021b. Neural bellman-ford networks: A general graph neural network framework for link prediction. Advances in Neural Information Processing Systems 34 (2021), 29476–29490.
@@
| Runtime (mins) | Offline | ||||||
| Baseline setup | 11h 24m | 2h 31m | 2h 29m | 50m | 4h 4m | 1h 30m | 0.820 |
| Half training GPU | 14h 7m | - | - | - | 6h 47m | - | - |
| Half inference CPU | 11h 43m | - | - | - | - | 1h 49m | - |
| Early stop patience * 2 | 19h 1m | - | - | - | 11h 41m | - | 0.858 |
| Remove edge features | 9h 27m | 2h 1m | 2h 25m | 46m | 2h 47m | 1h 28m | 0.806 |
@@ Appendix A Dataset Details
151515https://ogb.stanford.edu/docs/lsc/mag240m/. In order to evaluate the same link prediction task as we did for the internal datasets, the train/validation/test sets are generated by our Split Generator following a standard transductive link prediction split.
@@ B.1. Varying Level of Parallelism and Configurations
Half training GPU. We reduce the number of training GPUs from 16 to 8. This affects training only, since inference is done with CPUs.
Half inference CPU. We reduce the number of inference CPUs by half. This only affects inferencer and not trainer or other components.
Earyly stop patience * 2. Double the early stop patience will result in better loss optimization and metric, but slower training time.
# of positive samples * 2. We double the positive labels sampled for each training sample. This increase Subgraph Sampler & trainer run time.
@@ Appendix C Config Examples
for GiGL pipeline on MAG240M dataset for the task of link prediction: https://github.com/snap-research/GiGL/blob/main/examples/MAG240M/task_config.yaml. Detailed implementation of GiGL end-to-end pipeline on MAG240M dataset for link prediction is also provided as an example in the GiGL repositoryhttps://github.com/snap-research/GiGL/tree/main/examples/MAG240M.
@@
used w/ task_config.yaml specified above: https://github.com/snap-research/GiGL/blob/main/examples/MAG240M/resource_config.yaml.

### v2

@@ Acknowledgements.
- Zhu et al. (2021b) Zhaocheng Zhu, Zuobai Zhang, Louis-Pascal Xhonneux, and Jian Tang. 2021b. Neural bellman-ford networks: A general graph neural network framework for link prediction. Advances in Neural Information Processing Systems (2021).
@@
| Runtime (mins) | Offline | ||||||
| Baseline setup | 11h 24m | 2h 31m | 2h 29m | 50m | 4h 4m | 1h 30m | 0.820 |
| Half training GPU | 14h 7m | - | - | - | 6h 47m | - | - |
| Half inference CPU | 11h 43m | - | - | - | - | 1h 49m | - |
| Early stop patience * 2 | 19h 1m | - | - | - | 11h 41m | - | 0.858 |
| Remove edge features | 9h 27m | 2h 1m | 2h 25m | 46m | 2h 47m | 1h 28m | 0.806 |
@@ Appendix A Dataset Details
161616https://ogb.stanford.edu/docs/lsc/mag240m/. In order to evaluate the same link prediction task as we did for the internal datasets, the train/validation/test sets are generated by our Split Generator following a standard transductive link prediction split. We leverage the original public academic graph dataset (Hu et al., 2021, 2020b) and cast it to a homogeneous graph for our experiments by: (i) using a 768 size zero vector as the feature vector for each author node. (ii) Combine the new hydrated author nodes and original paper nodes with 768 dim float features into a single node table. (iii) Add the node degree of each node as the 769th feature. (iv) Combine the paper-cites-paper and the author-writes-paper edge table to a single edge table.
@@ B.1. Varying Level of Parallelism and Configurations
Half training GPU. Reduce the # of training GPUs from 16 to 8. This affects training only, since inference is done with CPUs.
Half inference CPU. Reduce the # of inference CPUs by half. This only affects inferencer and not trainer or other components.
Early stop patience * 2. Double the patience will result in better loss optimization and metric, but slower training time.
# of positive samples * 2. Double the positive labels sampled for each training sample. This increases Subgraph Sampler & Trainer run time.
@@ Appendix C Config Examples
for GiGL pipeline on MAG240M dataset for the task of link prediction: https://github.com/Snapchat/GiGL/blob/main/examples/MAG240M/task_config.yaml. Detailed implementation of GiGL end-to-end pipeline on MAG240M dataset for link prediction is also provided as an example in the GiGL repositoryhttps://github.com/Snapchat/GiGL/tree/main/examples/MAG240M.
@@
used with task_config.yaml specified above: https://github.com/Snapchat/GiGL/blob/main/examples/MAG240M/resource_config.yaml.